# Effective Game-Theoretic Motion Planning via Nested Search

### v1

@@ 6 Experiments and Results
We provide an experimental evaluation of our game-theoretic approach in driving and racing scenarios. We begin this section with implementation details and experiment setup, along with a discussion of the kinodynamic graph construction. We then proceed to a qualitative evaluation showcasing the behavior of the approach. Then we evaluate the computational and scalability aspects of our approach. We conclude with experiments examining the effect of the mismatch between the NE solutions chosen by individual robots, and the effects of limited short-horizon NE solutions, that can be computed using previous bimatrix approaches [liniger2019noncooperative, tikna2023graphs], as compared to our own approach (which is more scalable in terms of horizon length).
@@ 6.1 Implementation details and problem setting
The single-robot graphs $G_{i}$ are built offline in Python using CasADi Opti [Andersson2019] for BVP solving (see more information below). GTNS is implemented in C++ on a LEMON [dezsHo2011lemon] graph library backbone, and is executed online. All simulations were run on a 64-bit laptop with an Intel Core i9-14900HX CPU (2.20 GHz) and 32 GB RAM, running Windows 11. The experiments did not rely on GPU acceleration.
For the robot model, we consider the second-order bicycle model [lavalle2006planning], which models nonholonomic steering while remaining lightweight enough for real-time planning. The robot has the states, controls, and dynamics, of the form
$\mathbf{x}=\begin{bmatrix}x\\ y\\ \theta\\ v\\ \delta\end{bmatrix},\quad\mathbf{u}=\begin{bmatrix}a\\ \omega\end{bmatrix},\quad\dot{\mathbf{x}}=\begin{bmatrix}v\cos\theta\\ v\sin\theta\\ \tfrac{v}{L}\tan\delta\\ a\\ \omega\end{bmatrix}$,
respectively. Here $x$, $y$ and $\theta$ are the 2-D positions and heading of the car, $v$, and $\delta$ are its speed and steering angle, and $a$ and $\omega$ control their rate of change. $L$ is the distance between front and rear axles (wheelbase). We use a scale–model car of length $l=0.70\,\mathrm{m}$, width $w=0.20\,\mathrm{m}$, and wheelbase $L=0.50\,\mathrm{m}$. Collision checking uses a circular footprint of radius $r_{\mathrm{coll}}=\sqrt{l^{2}+w^{2}}/3$. See Appendix 0.A, Fig. 8 for visualization.
Kinodynamic graphs. For the single–robot graphs $G^{i}$, we consider two constructions. The grid graph (Fig. 2) samples a spatial lattice in $(x,y)$ (with $\Delta x=\Delta y=1\,\mathrm{m}$) and enumerates headings, speeds, and steering angles on discrete sets $(\theta,v,\delta)$ (with $\Delta v=1\,\mathrm{m/s}$ and $\Delta\theta,\Delta\delta$ specified per map). From each node we attempt directed connections to nearby nodes within a $k$-hop neighborhood of the lattice. The track graph (Appendix 0.A, Fig. 9) samples along a track centerline with lateral offsets: at each longitudinal sample we define a wayline (normal to the centerline), set the heading from the centerline tangent, and discretize $(v,\delta)$ as above. From each wayline $\ell$ we attempt connections from all nodes on $\ell$ to all nodes on any wayline $\ell^{\prime}$ with $|\ell^{\prime}-\ell|\leq k$. In both graph types, nodes represent states $(x,y,\theta,v,\delta)$ within bounds $\theta\in(-\pi,\pi]$, $v\in[0,10]\,\mathrm{m/s}$, $\delta\in[-\pi/2,\pi/2]$. Unless otherwise stated, robots share identical dynamics and graph $G^{i}$.
Our choice of graph type is guided by two complementary paradigms. The grid graph is task-agnostic: it covers arbitrary workspaces with obstacles and is suitable for general multi-robot scenes. By contrast, the track graph leverages road/track topology—known centerline and lane structure—to align headings with geometry, reduce unnecessary branching, and permit longer, better-conditioned local connections between nearby waylines. In practice, the track graph tends to be smaller and sparser, enabling faster search for driving/racing scenarios, whereas the grid graph provides uniform coverage when no privileged centerline exists.
Graph size is determined by the sampling resolutions $(\Delta x,\Delta y,\Delta\theta,\Delta v,\Delta\delta)$ and the connection parameter $k$. We choose these to meet a computational budget while preserving sufficient fidelity for the maneuvers of interest. Concretely, we set spatial and angular resolutions and $k$ to reach a desired ballpark of nodes/edges, then prune infeasible nodes/edges via kinodynamic feasibility and obstacle inflation. The resulting sizes per graph for a selection of graphs are reported in Table 1.
@@
$z_{j+1}=z_{j}+\Delta\tau\,f(z_{j},u_{j}),\qquad j=0,\ldots,N-1.$
In practice, this means we choose the piecewise-constant control sequence $u_{0},\ldots,u_{21}$ (and the corresponding knot-point states) that drives the model from the source to the target over $1\,\mathrm{s}$ while respecting bounds, with boundary conditions $z_{0}=\text{source},\ z_{N}=\text{target}$ and hard input limits $|a|\leq 5\,\mathrm{m/s^{2}}$, $|\omega|\leq 2\,\mathrm{rad/s}$. If the BVP is feasible, we add the corresponding directed edge and store its continuous trajectory. Edges are discarded if their continuous trajectory intersects obstacles inflated by $r_{\mathrm{coll}}$. Unless otherwise specified, each edge has unit cost.
Search heuristics For each robot $i$ we precompute all-pairs shortest-path distances on $G^{i}$ (APSP) and use $H(x)=\sum_{i}H^{i}(x^{i})$ where $H^{i}$ is the single-robot shortest distance from $x^{i}$ to $\mathcal{X}^{i}_{\mathrm{goal}}$. Because single‑robot distances satisfy the triangle inequality and joint costs add across robots, $H$ is consistent (and hence admissible) on the tensor‑product graph $G$. Moreover, during inner A* searches for robot $i$ on $G^{i}$ with dynamic obstacles induced by a fixed plan $\pi^{-i}$, $H^{i}$ remains a valid lower bound on $i$’s cost‑to‑go: the added constraints only restrict feasible motion and cannot reduce the optimal single‑robot cost.
@@ 6.2 Qualitative analysis
We present a set of scenarios mimicking interactions between multiple agents in autonomous driving and racing and discuss the results obtained using our algorithm.
Across tasks we vary two interpretable knobs that shape the resulting NE.
  1. 1.
    Priority in the global objective. The global objective aggregates the individual costs $J(\pi_{n})=\sum_{i=1}^{m}\alpha_{i}\,J^{i}(\pi^{i}_{n},\pi^{-i}_{n})$ to encode priorities via the weights $\alpha_{i}$.
  2. 2.
    Proximity (safety) inside each individual cost. Each $J^{i}$ includes a state-based penalty that discourages small separations from other robots:
    $c^{i}_{\mathrm{prox}}\!\left(\mathbf{x}^{i},\mathbf{x}^{-i}\right)\;=\;\frac{\lambda_{\mathrm{prox}}}{\max\!\left(\min_{j\in-i}\,\bigl\|P(\mathbf{x}^{i})-P(\mathbf{x}^{j})\bigr\|_{2},\ \varepsilon\right)},$
    where $\lambda_{\mathrm{prox}}\!\geq 0$, $P:\mathcal{X}\to\mathbb{R}^{2}$ returns the planar position $(x,y)$ of a state and $\varepsilon{=}10^{-3}$ prevents singularities (yielding a bounded, well-behaved, Lipschitz stage cost). Larger $\lambda_{\mathrm{prox}}$ promotes greater standoff.
In our experiments, each $J^{i}$ is the sum over $n$ edges of $\pi^{i}_{n}$ of unit edge cost plus the proximity term for each of the $n+1$ nodes along the way:
$J^{i}(\pi^{i}_{n},\pi^{-i}_{n})=n+\sum_{\mathbf{x}^{i}\in\pi^{i}_{n}}\!\Bigl(c^{i}_{\mathrm{prox}}(\mathbf{x}^{i},\mathbf{x}^{-i})\Bigr),$
so $\{\alpha_{i}\}$ weights between agents at the global level, while $\lambda_{\mathrm{prox}}$ shapes within-agent costs that the search optimizes on each edge.
Next, we discuss the scenarios and the solution obtained by GTNS.
Four-way intersection. We consider a single-lane four-way intersection with perpendicular approaches by two robots, objective $J=\alpha J^{1}+(1-\alpha)J^{2}$, $\alpha\in[0,1]$, $\lambda_{\mathrm{prox}}=0$. Fig. 3 shows that increasing $\alpha$ selects the NE in which Robot 1 proceeds and Robot 2 yields; decreasing $\alpha$ selects the opposite NE. This result illustrates the ability of our NE planner to select a safe NE in an explainable manner (via tuning of the objective function $J$), without requiring clipping of the dynamics or costs, or initial guesses (as commonly considered in previous work).
(a) $\alpha<0.5$
(b) $\alpha\geq 0.5$
@@
Racetrack overtake. We consider a S-curve racetrack segment to showcase priority-driven overtaking decisions between two robots. This setting probes how a scalar priority $\alpha$ biases raceline choice and pass timing within realistic behavioural dynamics trajectories captured by the NE framework. Here, we set $J=\alpha J^{1}+(1-\alpha)J^{2}$, $\alpha\in[0,1]$, $\lambda_{\mathrm{prox}}=0$.
Fig. 4 shows that the tilt of robot priority determines which robot targets the inside line at the final turn, and ultimately wins the race. It is important to note however, that even at high $\alpha$, Robot 1 is blocked at the early corner— illustrating that even in the best NE for Robot 1, overtaking early is not feasible due to kinodynamic/safety constraints. This too is a significant result as we can (using our tuned NE, and finding the best NE for a given robot) robustly express the lower bound on the cost of a given robot’s solution (within the space of all NEs). That is to say, that using our algorithm, we can discern, for example, when a loss is imminent, no matter what, and act accordingly (prioritize low damage and fuel burn, for example, conserving resources for subsequent tasks).
(a) $\alpha<0.5$
(b) $\alpha\geq 0.5$
@@
Following distance. Here, we consdier a more complex racetrack segment and evaluate how proximity weight affects behaviour in semi-collaborative setting, such as two cars on the same F1 team, with social welfare cost $J=\sum_{i}J^{i}$ and $\lambda_{\mathrm{prox}}\in[0,1]$.
The effect of the proximity penalty is clear in Fig. 5, with small $\lambda_{\mathrm{prox}}$ yielding close drafting, and larger values increasing the time-gap as the trailing car deliberately backs off (visible at the first corner). This result shows the strength of our algorithm for solving global NE, with (almost) generic costs.
(a) $\lambda_{\mathrm{prox}}=0.00$
(b) $\lambda_{\mathrm{prox}}=0.25$
(c) $\lambda_{\mathrm{prox}}\geq 0.50$
Figure 5: Following distance. Robot 1 (blue) trails at various distances behind Robot 2 (orange) as proximity is more heavily penalized.
Multi-lane highway merge. We consider a three-lane highway merge with three robots to explore homotopy classes of merge options as $\alpha$ and $\lambda_{\mathrm{prox}}$ trade priority against risk. We set $J=(1-\alpha)J^{1}+\alpha(J^{2}+J^{3})$, $\alpha\in[0,1]$, and $\lambda_{\mathrm{prox}}\in[0,2.5]$.
We can see in Fig. 1 complex behavioural changes. As priority for Robot 1 decreases and $\lambda_{\mathrm{prox}}$ increases, the NE transitions through: (i) aggressive zip-merge (tight gaps); (ii) calm zip-merge interleaving; (iii) yield-then-merge behind; (iv) over-cautious merge with enlarged headways. This shows the multiplicity of homotopy classes that often (if not always) exist in complex multi-robot systems, that are often impossible to capture with local NE solvers. Local methods can be sensitive to initialization and may remain within a single homotopy class; our search, by construction, explores all classes.
Opposing-lane overtake. We consider a two-lane bidirectional road with oncoming traffic and three robots to simulate risk-aware passing with heterogeneous capabilities ($G^{1}$ has $v\in[0,10],\ k=5$, $G^{2},G^{3}$ have $v\in[0,4],\ k=3$). We also test how $\alpha$ and $\lambda_{\mathrm{prox}}$ modulate dangerous overtake decisions and select among oncoming-lane vs. in-lane homotopy classes. We set $J=(1-\alpha)J^{1}+\alpha(J^{2}+J^{3})$, $\alpha\in[0,1]$, and $\lambda_{\mathrm{prox}}\in[0,1]$.
In Fig. 6 we see that with high priority and low $\lambda_{\mathrm{prox}}$, Robot 1 overtakes at the first viable opportunity. Increasing $\lambda_{\mathrm{prox}}$ and lowering priority, however, delays the pass until oncoming traffic clears, and beyond a certain threshold, the NE forgoes the overtake entirely. This too shows extreme behavioural changes across the different homotopy classes, as well as providing (as above) an upper/lower bound on potentially reachable behaviors/outcomes within the NE space.
(a) $\alpha=0.00,\ \lambda_{\mathrm{prox}}=0.00$
(b) $\alpha=0.25,\ \lambda_{\mathrm{prox}}=0.25$
(c) $\alpha\geq 0.50,\ \lambda_{\mathrm{prox}}\geq 0.50$
Figure 6: Opposing-lane overtake. Robot 1 (blue) tries to overtake Robot 2 (orange) via Robot 3’s (green) opposing lane. Different NE overtake strategies chosen as priority and proximity penalty modulated
@@ 6.3 Running time evalution
Table 1: Selected graphs with sizes and per-distance solve times. Each row corresponds to a map, number of robots ($m$). The rightmost block lists time(s) in seconds with the associated goal distance $D$ in parentheses, ordered by increasing $D$.
@@
66 0.363(8)
77 2.32(8)
Table 1 reports wall–clock solve times for GTNS across several maps, goal distances and number of robots. Dimensions are $|x|\times|y|\times|\theta|\times|v|\times|\delta|$, where $|\cdot|$ is the number of discrete values used for each state variable.
The Generic Free map is an obstacle–free workspace used to stress test expansion (maximal branching), whereas Generic Obstacles uses the same spatial dimensions but with randomly placed obstacles that prune many nodes/edges (lower branching).
@@
Goal distance $D$ is a conservative proxy for planning horizon $T$ as a time cap $T$ restricts search to the subgraph reachable within $\lfloor T/\Delta t\rfloor$ edges from the start; by instead varying the goal distance $D$, we probe increasingly far targets without imposing this cap.
We note that increasing $D$ does not substantially increase running time—a considerable strength of our approach: unlike optimization-based approaches—where running time typically scales with horizon due to the growing number of decision variables—our graph-based formulation keeps per-step local problems fixed-size and searches over a discrete action set, so increasing $D$ does not balloon the optimizer.
@@
By contrast, running time generally increases with the number of robots. The joint neighbor set grows combinatorially, interaction geometry becomes richer (more constraints and blocking situations), and the equilibrium check isNashEquilibrium triggers more single–robot A* calls (a principal computational bottleneck) as $m$ grows. An exception is the Crossroads map: its branching factor is small and interactions are simple, so adding robots does not markedly increase search effort.
@@
Across all settings, instance difficulty (map layout and interaction geometry) dominates running time. Tight merges, opposing–lane passes, and narrow corridors can make some two–robot cases slower—or even infeasible at a given discrete horizon—than multi–robot cases on simpler maps with few viable options. This can also explain why Generic Obstacles is faster on average than Generic Free: moderate obstacle density reduces branching without making the graph too sparse.
6.4 Comparison with previous short-horizon methods
In this section, we provide a comparison with previous work that can be most directly simulated in our graph-based framework. Specifically, we consider approaches that compute NE through an explicit enumeration of all possible solutions within a bimatrix formulation [liniger2019noncooperative, tikna2023graphs]. As reported in [liniger2019noncooperative], solving just $N=3$ prediction steps—which corresponds to an horizon of three steps in our setting—requires constructing and brute-force solving two payoff matrices with sizes on the order of $10^{12}$—an intractable task. To assess this limitation, we compare the solution cost obtained by GTNS using a long horizon of $17$ steps, with that obtained by GTNS attempting the same problem, but only invoked to solve subproblems with horizon $2$ in MPC fashion.
As expected, across multiple track scenarios (e.g. those in Fig. 4 and  5), both the individual costs $J^{i}$ and the team cost $J$ degraded relative to the long horizon. The behavior reflects myopia: short horizons optimize locally and replan, while the long horizon equilibrium, enabled by GTNS, anticipates future interactions and coordinates earlier. Although absolute numbers depend on discretization, the trend was consistent: long-horizon joint NE planning dominated short-horizon patching in both cost and qualitative safety. Costs in the short-horizon setting were often 2x the costs of their long-horizon counterparts.
6.5 Effect of NE mismatch
In the majority of tested cases, when agents executed in an open-loop fashion trajectories belonging to different NEs, collisions between robots were encountered. This underscores the need for a coordinated, team-wide equilibrium in behavior-aware motion planning. To address the concern that open-loop per-agent NEs are “obviously brittle,” we also evaluated an MPC variant: at every step each agent re-solved its own NE instance, where $J=J_{i}$, i.e., the global cost strictly optimizes the solution of the specific robot executing its trajectory from this NE. At every step, the robot resolved the problem towards the goal using this objective while observing the updated locations of the other robots, executed its step, and repeated the process. Such an approach is hoped to be more resilient as it could provide more opportunities for the robots to avoid colliding with each other. Unfortunately, even under this closed-loop regime, the lack of shared equilibrium selection caused systematic issues and in almost all scenarios (e.g. those in Fig. 6 and  3) a collision. In rare “best-case” scenarios, when interaction geometry was simple, (e.g. on the track in Fig. 5 with large start distance) mismatched individual NEs yielded similar solutions to those returned by the global planner.

### v2

@@ 6 Experiments and Results
We provide an experimental evaluation of our game-theoretic approach in driving and racing scenarios. We begin with implementation details and experiment setup, and then proceed to a qualitative evaluation showcasing the behaviors of the approach. Next, we evaluate the computational aspects of our approach and conclude with a comparison with LaValle-Hutchinson [lavalle2002optimal]. An additional experiment, which studies the effect of NE mismatch, is found in App. 0.B.1.
@@ 6.1 Implementation details and problem setting
The single-robot graphs $G_{i}$ are built offline in Python using CasADi Opti [Andersson2019] for BVP solving (see more information below). GTNS is implemented in C++ using the LEMON [dezsHo2011lemon] graph library. Simulations were run on a 64-bit Win11 laptop with an Intel Core i9-14900HX CPU (2.20 GHz) and 32 GB RAM.
For the robot model, we consider the second-order bicycle model [lavalle2006planning]. The robot has the states, controls, and dynamics, of the form
$\small\mathbf{x}=\begin{bmatrix}x\\ y\\ \theta\\ v\\ \delta\end{bmatrix},\quad\mathbf{u}=\begin{bmatrix}a\\ \omega\end{bmatrix},\quad\dot{\mathbf{x}}=\begin{bmatrix}v\cos\theta\\ v\sin\theta\\ \tfrac{v}{L}\tan\delta\\ a\\ \omega\end{bmatrix}$,}
respectively. Here, $x$, $y$ and $\theta$ are the 2-D positions and heading of the car, $v$, and $\delta$ are its speed and steering angle, and $a$ and $\omega$ control their rate of change. $L$ is the distance between front and rear axles (see visualization in App. 0.B, Fig. 8).
Kinodynamic graphs. For the single-robot graphs $G^{i}$, we consider two constructions. The grid graph (Fig. 2), which is task-agnostic, samples a spatial lattice in $(x,y)$ (with $\Delta x{=}\Delta y{=}1\,\mathrm{m}$) and expands said lattice with headings, speeds, and steering angles on discrete sets $(\theta,v,\delta)$. From each node, we attempt directed connections to nearby nodes within a $k$-hop neighborhood of the lattice. The track graph leverages the environment’s topology and samples along a track centerline with lateral offsets (App. 0.B, Fig. 7). Unless otherwise stated, robots share identical dynamics and graph $G^{i}$.
Graph size is determined by the sampling resolutions $(\Delta x,\Delta y,\Delta\theta,\Delta v,\Delta\delta)$ and the connection parameter $k$. We choose these to meet a computational budget while preserving sufficient fidelity for the maneuvers of interest. Concretely, we set spatial and angular resolutions and $k$ to reach a desired ballpark of nodes/edges, then prune infeasible nodes/edges via kinodynamic feasibility and obstacle inflation. The resulting graph sizes are reported in Table 1.
@@
Search heuristics. As the inner search isNashEquilibrium may be called on any $\pi_{x_{\textup{cand}}}$, we require an admissable heuristic to every node $x$, and not just to $\mathcal{X}_{\textup{goal}}$, thus for each robot $i$ we precompute all-pairs shortest-path (APSP) distances on $G^{i}$ and use $H(x)=\sum_{i}H^{i}(x^{i})$ (or the weighted equivalent when $J(\cdot)$ is weighted, see below) where $H^{i}$ is the single-robot shortest distance from $x^{i}$ to $\mathcal{X}^{i}_{\mathrm{goal}}$. Because single‑robot distances satisfy the triangle inequality and joint costs add across robots, $H$ is consistent (and hence admissible) on the tensor‑product graph $G$. Moreover, during inner A* searches for robot $i$ on $G^{i}$ with dynamic obstacles induced by a fixed plan $\pi^{-i}$, $H^{i}$ remains a valid lower bound on $i$’s cost‑to‑go: the added constraints only restrict feasible motion and cannot reduce the optimal single‑robot cost.
@@ 6.2 Qualitative analysis
We present a set of multi-agent scenarios mimicking interactions in autonomous driving and racing and discuss the results obtained by our algorithm. These experiments showcase the generality of our approach, and demonstrate the diversity of equilibria available, which are hard to capture with previous optimization-based methods.
Across tasks we vary two interpretable knobs that shape the resulting NE: The global objective aggregates the individual costs $J(\pi_{n})=\sum_{i=1}^{m}\alpha_{i}\,J^{i}(\pi^{i}_{n},\pi^{-i}_{n})$ to encode priorities via the weights $\alpha_{i}\in[0,1]$. Additionally, each $J^{i}$ includes a state-based penalty that encourages separation from other robots:
$c^{i}_{\mathrm{prox}}\!\left(\mathbf{x}^{i},\mathbf{x}^{-i}\right)\;=\;\frac{\lambda_{\mathrm{prox}}}{\max\!\left(\min_{j\in-i}\,\bigl\|\textup{proj}(\mathbf{x}^{i})-\textup{proj}(\mathbf{x}^{j})\bigr\|_{2},\ \varepsilon\right)},$
where $\lambda_{\mathrm{prox}}\!\geq 0$, $\textup{proj}:\mathcal{X}\to\mathbb{R}^{2}$ returns the planar position $(x,y)$ of a state and $\varepsilon{=}10^{-3}$ prevents singularities (yielding a bounded, well-behaved, Lipschitz stage cost). Larger $\lambda_{\mathrm{prox}}$ promotes greater standoff. In our experiments, each $J^{i}$ is the sum over $n$ edges of $\pi^{i}_{n}$ of unit edge cost plus the proximity term for each of the $n+1$ nodes along the way:
$J^{i}(\pi^{i}_{n},\pi^{-i}_{n})=n+\sum_{\mathbf{x}^{i},\mathbf{x}^{-i}\in(\pi^{i}_{n},\pi^{-i}_{n})}\!\Bigl(c^{i}_{\mathrm{prox}}(\mathbf{x}^{i},\mathbf{x}^{-i})\Bigr).$
Next, we discuss the scenarios and the solution obtained by GTNS. An additional experiment that showcases the effect of $\lambda_{\mathrm{prox}}$ on following distance in a semi-collaborative setting, can be found in App. 0.B, Fig. 5.
@@
Four-way intersection. We consider a single-lane four-way intersection with perpendicular approaches by two robots, objective $J=\alpha J^{1}+(1-\alpha)J^{2}$, $\alpha\in[0,1]$, $\lambda_{\mathrm{prox}}=0$. For $\alpha\geq 0.5$ the NE selected by our planner has Robot 1 (blue) crossing from left to right while Robot 2 (orange) yields and waits for the intersection to clear before proceeding from bottom to top. For $\alpha<0.5$, the NE is symmetric: Robot 2 crosses first and Robot 1 yields. This result illustrates the ability of our NE planner to select a safe NE in an explainable manner (via tuning of the objective function $J$), without requiring clipping of the dynamics or costs, or initial guesses (as commonly considered in previous work).
Racetrack overtake. We consider a racetrack segment to showcase priority-driven overtaking decisions between two robots. This setting probes how a scalar priority $\alpha$ biases raceline choice and pass timing within realistic behavioral dynamics trajectories. We set $J=\alpha J^{1}+(1-\alpha)J^{2}$, $\alpha\in[0,1]$, $\lambda_{\mathrm{prox}}=0$. Fig. 3 shows that the tilt of robot priority determines which robot targets the inside line at the final turn, and ultimately wins the race. Note that even at high $\alpha$, Robot 1 is blocked at the early corner—illustrating that even in the best NE for Robot 1, overtaking early is not feasible due to kinodynamic/safety constraints. Thus, via a tuned NE our algorithm can robustly express the lower bound on the cost of a given robot’s solution (within the space of all NEs).
[…]
@@
Multi-lane highway merge. We consider a three-lane highway merge with three robots (Fig. 1) to explore homotopy classes of merge options as $\alpha$ and $\lambda_{\mathrm{prox}}$ trade priority against risk. We set $J=(1-\alpha)J^{1}+\alpha(J^{2}+J^{3})$, $\alpha\in[0,1]$, and $\lambda_{\mathrm{prox}}\in[0,2.5]$. As priority for Robot 1 decreases and $\lambda_{\mathrm{prox}}$ increases, the NE transitions through distinct maneuvers. This shows the multiplicity of homotopy classes that often (if not always) exist in complex interactions, that are often impossible to capture with local NE solvers.
Opposing-lane overtake. We consider a two-lane bidirectional road with oncoming traffic and three robots to simulate risk-aware passing with heterogeneous capabilities ($G^{1}$ has $v\in[0,10],\ k=5$, $G^{2},G^{3}$ have $v\in[0,4],\ k=3$). We also test how $\alpha$ and $\lambda_{\mathrm{prox}}$ modulate dangerous overtake decisions and select among oncoming-lane vs. in-lane homotopy classes. We set $J=(1-\alpha)J^{1}+\alpha(J^{2}+J^{3})$, $\alpha\in[0,1]$, and $\lambda_{\mathrm{prox}}\in[0,1]$. In Fig. 4 we see that with high priority and low $\lambda_{\mathrm{prox}}$, Robot 1 overtakes at the first viable opportunity. Increasing $\lambda_{\mathrm{prox}}$ and lowering priority, however, delays the pass until oncoming traffic clears, and beyond a certain threshold, the NE forgoes the overtake entirely. This shows behavioral changes across the different homotopy classes.
[…]
Figure 4: Opposing-lane overtake. Robot 1 (blue) tries to overtake Robot 2 (orange) via Robot 3’s (green) opposing lane. Different NE overtake strategies chosen as priority and proximity penalty are modulated.
@@ 6.3 Running time evalution
Table 1: Running time evaluation for GTNS. Each row corresponds to a map, number of robots ($m$). The rightmost block lists time(s) in seconds with the associated goal distance $D$ in parentheses. The Generic Free map is an obstacle-free workspace used to stress test expansion (maximal branching). Generic Obstacles uses the same spatial dimensions but has random obstacles, pruning many nodes/edges (lower branching).
@@
Table 1 reports wall-clock solve times (not including preprocess time for graph generation) for GTNS across several maps, goal distances and number of robots. Goal distance $D$ is a proxy for planning horizon $T$ as a time cap $T$ restricts search to the subgraph reachable within $\lfloor T/\Delta t\rfloor$ edges from the start; by instead varying the goal distance $D$, we probe increasingly far targets without imposing this cap.
@@
We note that increasing $D$ does not substantially increase running time—a considerable strength of our approach. Unlike optimization-based approaches—where running time typically scales with horizon due to the growing number of decision variables—our graph-based formulation keeps per-step local problems fixed-size and searches over a discrete action set, so increasing $D$ does not balloon the optimizer. By contrast, running time generally increases with the number of robots. The joint neighbor set grows combinatorially, interaction geometry becomes richer (more constraints and blocking situations), and the equilibrium check isNashEquilibrium triggers more single-robot A* calls (a principal computational bottleneck) as $m$ grows. An exception is the Crossroads map: its branching factor is small and interactions are simple, so adding robots does not markedly increase search effort.
@@
Across all settings, instance difficulty (map layout and interaction geometry) dominates running time. Tight merges, opposing-lane passes, and narrow corridors can make some two-robot cases slower—or even infeasible at a given discrete horizon—than multi–robot cases on simpler maps with few viable options. This can also explain why Generic Obstacles is faster on average than Generic Free: moderate obstacle density reduces branching without making the graph too sparse. See App. 0.B, Fig. 9 for visualization.
6.4 Comparison to LaValle-Hutchinson method
As noted in Sec. 1, the LaValle-Hutchinson method [lavalle2002optimal] also leverages graph search techniques and yields equilibrium guarantees. Here, we study the computational burden of generating, pairwise verifying, and explicitly maintaining all non-dominated candidate trajectories (up to one representative per equivalence class) within that approach. Unsurprisingly, in maps with low branching, and short horizons (e.g., Crossroads, Track), and two robots, the two algorithms are almost comparable: our method achieves a modest speedup of roughly $1.5$-$2\times$. For dense graphs with high branching factors and long horizons (e.g., Two-Lane-Road, Generic Free), however, the speedup reaches 22-33 orders of magnitude.
The main reason is that the LaValle-Hutchinson algorithm must carry every non-dominated partial trajectory (equivalency class) at each joint state. As the branching factor and horizon grow, the number of distinct cost vectors at a single node can grow combinatorially, so each additional layer multiplies both the number of wavefront paths and the cost of dominance checks. In contrast, our planner explores the composite state space with a best-first strategy guided by single-robot distance heuristics and aggressive dominance, so it (a) explores more promising areas first (only) and (b) effectively prunes entire sub-trees. Thus, its complexity scales primarily with the number of joint states actually visited, rather than with the total number of distinct joint trajectories, which makes high-branching and large-horizon regimes much easier to handle in practice.